OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)?

OpenStudy (anonymous):

m=6/-2=-3

OpenStudy (anonymous):

b=5

OpenStudy (anonymous):

these are the choices: y – 11 = –3(x + 2) y – 11 = –3(x – 2) y – 5 = 3(x + 2) y – 5 = –3(x + 2)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Thank you!